The perimeter of a rectangle is 208 inches. The length exceeds the width by 52 inches. Find the length and the width.
Answer by oberobic(2304)

P = 2*(L+W) = 208
.
L+W = 208/2 = 104
.
L = 104-W
.
L = W+52
.
L=L, so
104-W = W+52
52 = 2W
W = 26
.
L = W+52 = 26+52 = 78
.
Check answer with perimeter.
L+W = 26+78 = 104
P = 2*(L+W) = 2*104 = 208
Correct.
.
Answer: L=78, W=26
.
Done.